# Automated Review Generation Method Based on Large Language Models

### v1

@@ 4.5 Data mining
Code Availability
The custom code developed for this research is central to our conclusions and is made available to ensure transparency and reproducibility of our results. The codebase, including all relevant custom scripts and mathematical algorithms, has been open-sourced under the Apache 2.0 license and is accessible via our GitHub repository at [https://github.com/TJU-ECAT-AI/AutomaticReviewGeneration]. We encourage users to review the license for any usage restrictions that may apply. As stated in the text, all LLMs invoked in this article are Claude2.
It is important to note that our published graphical user interface (GUI) leverages certain APIs for functionality, which, due to legal and regulatory requirements, necessitate that users provide their own API keys. This requirement is detailed in the documentation accompanying the code repository to assist users in setting up and utilizing the GUI effectively.
Acknowledgments
The authors acknowledge the financial support from the National Key R&D Program of China (2022YFE0102000), the National Natural Science Foundation of China (No. 22121004, U22A20409), the Haihe Laboratory of Sustainable Chemical Transformations, the Program of Introducing Talents of Discipline to Universities (BP0618007) and the XPLORER PRIZE. The authors also appreciate the generous computing resources provided by the High Performance Computing Center of Tianjin University.
Author contributions
J.G and Z.Z. conceived and supervised the project. S.W. and X.M. designed the research. S.W. led the manuscript drafting. D.L. and L.L. validated the scientific reliability of the method. X.S. and X.C. innovated at the intersection of computational models and catalytic science. X.L., R.L., and C.P. conducted data analysis and validation. All authors wrote and revised this manuscript.
Competing interests

### v3

@@ 7 DATA AVAILABILITY
8 CODE AVAILABILITY
The custom code developed for this research is central to our conclusions and is made available to ensure transparency and reproducibility of our results. The codebase, including all relevant custom scripts and mathematical algorithms, has been open-sourced under the Apache 2.0 license and is accessible via our GitHub repository at [https://github.com/TJU-ECAT-AI/AutomaticReviewGeneration]. We encourage users to review the license for any usage restrictions that may apply.
As stated in the text, all LLMs invoked in this article are Claude 2, except for the Evaluation of generated review quality section which uses Claude 3.5 Sonnet, Qwen2-72b-Instruct, and Qwen2-7b-Instruct. It is important to note that our data processing work, except for the evaluation section, was completed prior to November 21, 2023, utilizing the then-latest available Claude 2 API version [https://www.anthropic.com/news/claude-2, https://www.anthropic.com/news/claude-2-1]. Anthropic did not publish specific minor version numbers within the Claude 2 series, only distinguishing between Claude 2, Claude 2.1, and the subsequent Claude 3 series. Our proposed framework demonstrates good adaptability, with overall effectiveness increasing as the performance of the underlying model improves. This characteristic has been amply demonstrated in our evaluation work, indicating that the framework’s efficacy is not strictly dependent on any particular model version. The use of different LLMs in the Evaluation of generated review quality section was primarily to assess the performance of the latest and most powerful open-source and closed-source models (as of September 2024) under the method described in this paper.
Our published graphical user interface (GUI) leverages certain APIs for functionality, which, due to legal and regulatory requirements, necessitate that users provide their own API keys. This requirement is detailed in the documentation accompanying the code repository to assist users in setting up and utilizing the GUI effectively.
9 CONFLICT OF INTEREST STATEMENT